OpenStudy (angela210793):

Some help pls...^_^ x^2-12x-2a=0 Find a=? so that X1=2X2 (x1,x2 are the roots of the eq.) (I guess it has smth to do with Vieta's formula which says X1+x2=-b/a and x1*x2=c/a

OpenStudy (anonymous):

-16

OpenStudy (angela210793):

O.o some explanation?

OpenStudy (anonymous):

check if its right

OpenStudy (anonymous):

if u use vieta s formula then the answer is -16 no a cannot be checked for all real numbers

OpenStudy (anonymous):

From the equation \(x_1+x_2=-\frac{b}{a}\), we have \(3x_2=12 \implies x_2=4\). Substitute that in the equation \(x_1 x_2=\frac{c}{a} \implies 2(x_2)^2=-2a \implies -a=4^2 \implies a=-16\).

OpenStudy (angela210793):

@vinitkocharekar This was the explanation I needed :P:P Thnx anyway :) @Anwar Thnx a lot ^_^

OpenStudy (anonymous):

You're welcome!